OpenStudy (anonymous):

I think this one is a bit different, so help solving for x? ._. (Posting the equation in comments)

OpenStudy (anonymous):

\[e ^{x}e ^{(x+1)}=1\]

OpenStudy (anonymous):

I know e goes with natural logarithms... need help isolating x though, since it's an exponent in this case...

OpenStudy (anonymous):

try multiplying both sides by the ln function

OpenStudy (anonymous):

well, ln1 is 0, and lne would cancel the e out.... I think, right?

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

So... x(x+1) = 0?

OpenStudy (anonymous):

you cant take the ln of zero so after you take the first ln to get rid of 1 e, you cant go any further so x would approach infinity. thats what i think but i may be wrong. do you know the answer?

OpenStudy (ranga):

\[\Large a^m * a^n = a^{m+n}\]

OpenStudy (anonymous):

Realize that \[e^x*e^(x+1) = e^(2x+1)\]

OpenStudy (anonymous):

So your equation is simplified to e^(2x+1) = 1

OpenStudy (anonymous):

Ohh okay.. so if I found the natural logarithms of both sides it would leave me with 2x + 1 = 0?

OpenStudy (anonymous):

bingo

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

So ultimately x = -1/2

OpenStudy (ranga):

yes.

OpenStudy (isaiah.feynman):

Oh no! I missed a nice exponential equation to solve!

OpenStudy (ranga):

you can put it back in the original equation and verify it fpr yourself.

OpenStudy (anonymous):

Yay! And awh Isaiah I'm sure I'll have more to come eventually xD I'll tag you in them if I have anymore questions.